Francisco has 7 quarters in his pocket of these 3 depict the state of Delaware 2 depict Georgia 1 depicts Texas and 1 depicts Pennsylvania. Francisco removes 1 quarter from his pocket and then removes a second quarter without replacing the first what is the probability that both will be Delaware quarters? Is this independent or dependent.
Answer by edjones(8007)

3/7 * 2/6 = 6/42 = 1/7 he probability that both will be Delaware quarters.
It is dependent because it would be different if a Georgia quarter was picked first.
.
Ed